OpenStudy (anonymous):

solve quadratic equation formula X2+9x+20=0

OpenStudy (anonymous):

x= -5, -4.. is that what your looking for?

OpenStudy (chihiroasleaf):

you can factorize the equation become (x+p)(x+q)=0 with p+q=9 and pq=20

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yes that helped thatnk you

OpenStudy (anonymous):

Np.